Percent of capital structure:
Debt 30 %
Preferred stock 15
Common equity 55
Additional information:
Bond coupon rate 10 %
Bond yield 8 %
Dividend, expected common $4.00
Dividend, preferred $11.00
Price, common $55.00
Price, preferred $118.00
Flotation cost, preferred $2.80
Corporate growth rate 7 %
Corporate tax rate 35 %

Calculate the weighted average cost of capital for Genex Corporation. Line up the calculations in the order shown in Table 11-1.
Answer #1
Cost of debt = Bond yield*(1-t) = 8%*(1-35%) = 5.20%
Cost of preferred stock = Preferred dividend/(Price-Flotation cost) = 11/(118-2.80) = 9.55%
Cost of equity per constant dividend growth formula = Next expected dividend/Price+Growth rate = 4/55+0.07 = 14.27%
WACC = Weight of debt*Cost of debt+Weight of preferred stock*Cost ofpreferred stock+Weight of equity*Cost of equity.
CALCULATIONS: After tax cost Weights WACC
Debt 5.20% 30% 1.56%
Preferred stock 9.55% 15% 1.43%
Common equity 14.27% 55% 7.85%
Weighted average cost of capital 10.84%